OpenStudy (anonymous):

evaluate the indefinite integra1 as an infinite series : int sin(x^2)dx

OpenStudy (anonymous):

\[\sin(x) = \sum_{0}^{\infty} \frac{ (-1)^{n} x^{2n+1}}{ (2n+1)! }\] sin(x^2) --> plug x = (x^2). find integral in the fashion of \[\int\limits_{}^{} x^{n} = \frac{ x^{n+1} }{ n+1 } +C\]

OpenStudy (anonymous):

∫x^n=(x^n+1)/(n+1)+C this is what you evaluated?

OpenStudy (anonymous):

how did you get that series so easily?

OpenStudy (anonymous):

wait wait, so i need to plug x with x^2?

OpenStudy (anonymous):

i know it by memory :P. it is not what i evaluated. that's the general rule for integrating function that look like power series (polynomials). and yes you plug x^2 everywhere there is x in the power series of sinx

OpenStudy (anonymous):

but there's only one x exists in the power series of sinx ?

OpenStudy (anonymous):

it sounds so simple now then???

OpenStudy (anonymous):

so same series but x^2^(2n+1) ???????????

OpenStudy (anonymous):

the proper way of doing it would be x^(2(2n+1))

OpenStudy (anonymous):

but yes. based on the same series

OpenStudy (anonymous):

was this that easy one???????

OpenStudy (anonymous):

idk what you mean

OpenStudy (anonymous):

this could be done easily if i knew power series of sinx???

OpenStudy (anonymous):

yes. that is the only way to do it. to base it on the power series of sinx. finding the power series of sin(x^2) right away would be very long

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

hi unkle

OpenStudy (anonymous):

hey unkle can you look over it? if you agree or not?

OpenStudy (unklerhaukus):

( i am trying a substitution method , ill post it when i get it to work )

OpenStudy (anonymous):

okay thank you! :)

OpenStudy (unklerhaukus):

*if

OpenStudy (unklerhaukus):

[i cannot help here]

OpenStudy (anonymous):

simpler than you'd think: sin(x^2) = \[\int\limits_{}^{} \sum_{0}^{\infty} \frac{ (-1)^{n} x^{2(2n+1)}}{ (2n + 1)! } = \int\limits_{}^{} \sum_{0}^{\infty} \frac{ (-1)^{n} x^{4n+2}}{ (2n + 1)! } \] =\[\sum_{0}^{\infty} \frac{ (-1)^{n} x^{4n+3}}{ (2n + 1)!*(4n +3)} \]

OpenStudy (anonymous):

first line of my last post should read integral of sin(x^2), not just sin(x^2)